# DesTest: A Decentralised Testing Architecture for Improving Data Accuracy of Blockchain Oracle

### v1

@@ Abstract
Blockchain technology ensures secure and trustworthy data flow between multiple participants on the chain, but interoperability of on-chain and off-chain data has always been a difficult problem that needs to be solved. To solve the problem that blockchain systems cannot access off-chain data, oracle is introduced. however, existing research mainly focuses on the consistency and integrity of data, but ignores the problem that oracle nodes may be externally attacked or provide false data for selfish motives, resulting in the unresolved problem of data accuracy. In this paper, we introduce a new decentralized testing architecture (DesTest) that aims to improve data accuracy. A blockchain oracle random secret testing mechanism is first proposed to enhance the monitoring and verification of nodes by introducing a dynamic anonymized question-verification committee. Based on this, a comprehensive evaluation incentive mechanism is designed to incentivize honest work performance by evaluating nodes based on their reputation scores. The simulation results show that we successfully reduced the discrete entropy value of the acquired data and the real value of the data by 61.4%.
Keywords:
Blockchain Oracle Random Test Incentive Mechanism.
1 Introduction
Blockchain is a system employing distributed ledger technology, characterized by features such as immutability, transparency, and decentralization [10]. Smart contracts, as a crucial component of blockchain technology, enable the execution of automated contracts and protocols on distributed ledgers, thereby providing greater programmability to blockchain systems. However, smart contracts face limitations in directly incorporating data from the external world, constraining their functionality and application scope.
Oracle networks, serving as an intermediary platform between blockchain and the external world, act as a trusted third party, bridging the information gap between the blockchain and the external world, thereby enhancing the resilience and practicality of the blockchain ecosystem [14]. The most significant challenge in incorporating external data into the blockchain is ensuring the accuracy and trustworthiness of the data, known as the ”oracle problem” [4]. There are currently some studies on how to solve this problem. For instance, TownCrier [18] had proposed a centralized solution based on a Trusted Execution Environment (TEE), utilizing Intel Software Guard Extension (SGX) to ensure the authenticity of responses from HTTPS queries. However, if the TC host server had been compromised, there would have been an increase in system energy consumption and a single point of failure. ChainLink [5] introduced a decentralized and distributed trust model that spanned on-chain and off-chain components, securely pushing data between smart contracts and oracles. Yet, on-chain aggregation might have incurred high gas fees. Taghavi et al.’s research [16] suggested a reinforcement learning-based method for selecting reliable and cost-effective oracles by assigning reputation values, making it unprofitable to report false data. Despite extensive research on the trustworthiness of oracles, challenges persist in the entire process of introducing data into the blockchain. Oracle networks remain vulnerable to attacks or provide deceptive data due to malicious intentions. The inclusion of such malicious data on the blockchain can erode trust in the entire blockchain ecosystem.
@@ 1 Introduction
In this paper, we introduce DesTest, a novel decentralized testing framework designed to improve data accuracy. DesTest features a random secret testing mechanism for blockchain oracle systems, revitalizing node monitoring and verification by using a dynamic, anonymized question-verification committee to safeguard data integrity and authenticity. Additionally, we develop an evaluation incentive mechanism that accounts for node reputation scores, offering a fair and transparent assessment of node performance. This approach encourages nodes towards greater honesty and efficiency, ultimately boosting data accuracy.
@@
  • We design a comprehensive evaluation incentive mechanism, based on the node’s reputation score for in-depth evaluation, high reputation value will increase the probability of becoming a working node and the rewards will be increased, aiming to motivate oracle nodes and improve the accuracy of data.
@@
  • We creatively propose the Decentralised Testing (DesTest) architecture and conduct a series of experimental evaluations. The experimental results show that the proposed scheme reduces the discrete entropy value of the acquired data with respect to the true value of the data by 61.4% compared to the existing baseline.
@@
The remaining sections of this paper are organized as follows. Section 2 discusses related work. Section  3 describes the main elements of the mechanism. Section 4 designs the incentive mechanism. Section 5 presents simulation results and analysis, and Section 6 provides the conclusion of this paper.
2 Related Work
@@
There is a large amount of research dedicated to solving the blockchain oracle problem, which falls into two main approaches: reputation-based and voting-based. These solutions provide trustworthy mechanisms when introducing external data into the blockchain.
2.1 Reputation-Based Oracle
@@
The data provided by oracle is safeguarded from tampering by designing different mechanisms based on reputation. DiOr-SGX [17] uses hardware protection provided by Intel Software Guard Extensions (SGX) to select reputable nodes as leaders for data transfer in a trusted and secure execution environment. Witnet [7] selected nodes to complete RAD tasks and obtain block-producing opportunities based on their reputation values, with higher reputation indicating greater responsibilities. Pasdar et al. [15] developed a livestock blockchain oracle (LBO) that selected nodes based on the oracle’s reputation/performance metrics, addressing the issue of accessing off-chain private sensitive data. Madine et al. [11] designed a reputation system to identify abnormal behavior in oracles, kicking out nodes with low scores from the oracle network.
2.2 Voting-Based Oracle
@@
Oracle nodes play roles such as voters and validators to cast votes. ASTRAEA [2] introduced multiple entity users, including validators, submitters, and voters. Submitters delivered challenges into the system, and voters adopted low-risk/low-reward strategies, while validators pursued high-risk/high-reward strategies. Deepthought [8] combined the voting system derived from ASTRAEA with user reputation to reward the most honest users and reduce corruption risks caused by adversarial users or lazy voters. Cai et al. [6] presented a scoring scheme based on peer prediction and nonlinear betting rules, aiming to extract subjective data truthfully. Gigli et al. [9] designed oracles that required consensus on specified functions and conducted voting.
@@
The design schemes for the aforementioned oracles primarily focus on the consistency and integrity of data but do not adequately consider the impact of data accuracy on the trustworthiness of oracles. Currently, there is almost no mention in existing literature of testing methods to assess the accurate feedback of data from oracle nodes. Oraichain [1] proposed quality testing for AI models, allowing AI providers to earn fees only if their AI models passed these tests, primarily incentivizing AI providers to improve the accuracy of AI models. It is mentioned in  [16] that in choosing the most valuable oracle, other oracles will test that oracle randomly, but will be vulnerable to targeted attacks or submission of spoofed data. Simultaneously, due to the public and transparent nature of blockchain, the testing process faces constraints, particularly making it challenging to evaluate the honesty of data feeding by oracle nodes.
3 Overview
To enhance the security and reliability of the oracle system, we have designed a new random secret testing mechanism. The key to this mechanism lies in the mixed publication of requests containing standard answers and regular requests, with decentralized oracles being responsible for executing this task. By introducing this hybrid testing approach, we can effectively conduct random and unpredictable checks on oracle nodes, thereby minimizing the potential for malicious behavior.
3.1 DesTest Architecture
@@
Considering the potential for dishonesty and unreliability among oracle nodes, leading to malicious nodes arbitrarily modifying data, creating confusion, or intentionally favoring one party for personal gain, we have designed and implemented the DesTest architecture, as shown in Figure 1. It is mainly divided into four modules: Smart Contract, Oracle Nodes, Data Sources, and Test Cases.
3.1.1 Smart Contract.
@@
Deploy user contracts and oracle contracts on the blockchain, with the former handling blockchain user-initiated data requests and the latter responding and providing feedback [13].
Registration Contract.
Becoming an oracle node by pledging virtual currency, and at the same time, registration contracts to record the pledged amount and payment address, provide strong protection against Sybil’s attacks.
Payment Contract.
@@
The user deposits fees into the contract as a payment mechanism for the oracle node. Upon successful completion of a task, the deposited amount is converted into a substantial reward for the oracle node, ensuring fair compensation or else punishment.
Reputation Contract.
@@
The system assigns an initial reputation value to each node. At the end of the task, the reputation contract updates the reputation value of the node.
Proxy Contract.
@@
Provides an on-chain interface for user contracts and defines the API interface for oracle nodes to fetch data off-chain.
3.1.2 Oracle Nodes.
@@
3.1.3 Data Sources.
The prophecy machine gets the task and accesses the information through API to off-chain data sources. These API interfaces are defined in the agent contract.
3.1.4 Test Cases.
Committee members obtain the examples needed for their test tasks from the test case repository [3] for distribution. The test case repository is one of the tools that provides committee members of the Prophecy Machine system with rich, reliable test data.
Figure 1: Overview of DesTest.
3.2 Workflow of Random Testing
Figure 2 illustrates the interactions between various modules.
3.2.1 Request for publication.
When a user needs to fetch data, they first submit a request $q$ to the blockchain through the smart contract (Step 1). The question-verification committee generates a test data request $Q$ at a certain moment and sends it to the on-chain smart contract (Steps 2.2 and 2.3). If the request is for on-chain data, the user contract directly fetches the blockchain data. If the request is for off-chain data, the oracle contract receives and responds to the request. After receiving the user contract’s request, the blockchain will invoke the oracle contract and pass relevant parameters to it (Step 3).
3.2.2 Node Selection.
Off-chain oracle nodes, upon detecting the request $Q$ or $q$, use a Verifiable Random Function (VRF) to select nodes for the question-verification committee (Step 2.1). Simultaneously, we design a reputation-based node selection scheme to choose trusted nodes for off-chain data-feeding (Step 4).
3.2.3 Data Fetch.
The selected decentralized oracle nodes fetch data from off-chain data sources, such as servers, cloud platforms, and the internet, using the defined API (Step 5).
3.2.4 Request Response.
Oracle nodes data-feeding into the smart contract on the blockchain (the node executing the test data request also feeds the fetched data into the blockchain). The oracle contract returns the response to the user contract and simultaneously sends the response to all other oracle nodes (Step 6).
3.2.5 Result Verification.
Committee members retrieve the data delivered by oracle nodes from the blockchain. When designing test cases, the case results are explicit. Committee members identify the oracle node that received the test request $Q$ and verify whether the data fetched by that node matches the case results, thus determining whether the oracle node exhibits malicious behavior. The committee informs the oracle contract of the results, and the contract updates the reputation values of oracle nodes accordingly (Step 7).
@@ 3.2 Workflow of Random Testing
3.3 Design of Random Testing
@@
3.4 Node Selection Strategy
@@
3.5 Structure of the question-verification committee
We designed three types of nodes for the question-verification committee, which are described in detail below.
3.5.1 Questioner.
The questioner’s task is to initiate a query of off-chain data. This selects a test case, publishes the task in the blockchain ecosystem, and then waits for the smart contract to execute the issued request $Q$.
3.5.2 Judge.
Judge upon the detection of malicious endeavors, the questioner is summarily excised from the proceedings and subjected to punitive measures, concomitant with the statement of new participants. In this paradigm, the selection of a substitute questioner is predicated on a randomized algorithm, ensuring procedural integrity and fairness. Should a validator discern and report malevolent conduct by the questioner, the judge is tasked with deliberating the veracity of such claims. A determination of guilt results in the accuser assuming the role of the questioner for the ensuing cycle, whereas exoneration entails penalization of the accuser, thus maintaining a self-regulating ecosystem [19].
3.5.3 Validator.
The validator is responsible for retrieving the blockchain and validating the data provided by oracle. The results of these validations are then transmitted to the reputation contract, which facilitates the calculation of the reputation score. This mechanism ensures continuous evaluation and validation of off-chain data, thus consolidating the reliability of the system.
@@ 3.5 Structure of the question-verification committee
4 Incentive Mechanism
4.1 Reputation of Data-Feeding Oracle
@@
We calculate the reputation score of a node based on its historical accumulated reputation $\mathcal{R}_{acci}$, response time $\mathcal{T}_{i}$, accuracy in completing random testing tasks $\mathcal{AC}$, and reputation weight $\mathcal{RW}$.
@@
4.2 The Reputation of Committee Members
@@ 4.3 Punishment Mechanism
If error $\text{If error instances}>\theta,\text{ then execute}\left\{\begin{array}[]{l}\text{Add to blacklist},\\ \text{Deduct all deposits},\\ \text{Remove node from oracle network}.\end{array}\right.$ (11)
5 Simulation Result
@@
a $s$ = 10%
b $s$ = 20%
c $s$ = 30%

### v2

# DecTest: A Decentralised Testing Architecture for Improving Data Accuracy of Blockchain Oracle

@@ Abstract
Blockchain technology ensures secure and trustworthy data flow between multiple participants on the chain, but interoperability of on-chain and off-chain data has always been a difficult problem that needs to be solved. To solve the problem that blockchain systems cannot access off-chain data, oracle is introduced. However, existing research mainly focuses on the consistency and integrity of data, but ignores the problem that oracle nodes may be externally attacked or provide false data for selfish motives, resulting in the unresolved problem of data accuracy. In this paper, we introduce a new Decentralized Testing architecture (DecTest) that aims to improve data accuracy. A blockchain oracle random secret testing mechanism is first proposed to enhance the monitoring and verification of nodes by introducing a dynamic anonymized question-verification committee. Based on this, a comprehensive evaluation incentive mechanism is designed to incentivize honest work performance by evaluating nodes based on their reputation scores. The simulation results show that we successfully reduced the discrete entropy value of the acquired data and the real value of the data by 61.4%.
I INTRODUCTION
Blockchain is a system that uses distributed ledger technology with features such as immutability, openness and transparency, and decentralisation [1]. This technology has already been applied in several fields such as the Internet of Things [2], healthcare data management [3], finance [4] and supply chain [5] and is constantly expanding into more fields. Smart contracts, as a crucial part of blockchain technology, make it possible to execute automated contracts and protocols on distributed ledgers, thus bringing greater programmability to the blockchain system [6]. However, the inability of smart contracts to bring in data directly from the outside world is limited by their functionality and application scope.
As an intermediary platform connecting the blockchain to the outside world, oracle acts as a trusted third party, bridging the information gap between the blockchain and the outside world and making the blockchain ecosystem more resilient and practical [7]. One of the biggest challenges in introducing external data into the blockchain is how to ensure the accuracy and trustworthiness of the data, which is the famous ”oracle problem” [8, 9].
Some research has been devoted to solving this problem. For example, TownCrier [10] designs a centralized solution based on the Trusted Execution Environment (TEE), which uses the Intel Software Guard Extension (SGX) to obtain data and verify the accuracy of the data through smart contracts. Once the server of the TC host is attacked, there will be a single point of failure. So ChainLink [11] proposed a decentralized and distributed trust model across on-chain and off-chain components to address the single point of failure, although using a lightweight on-chain aggregation mechanism to remove outliers in oracle data proposals, it is still not able to combat fraudulent or deceptive data. In addition, Taghavi et al. [12] proposed a method for selecting reliable and economical oracles using a reinforcement learning model that selects one or two oracles at a time for data acquisition, with the risk of targeted attacks.
Despite extensive research on the issue of data accuracy in oracle, the whole process of introducing data into the blockchain still faces the problem of oracle being attacked or providing false or deceptive data for selfish motives. With these malicious data on the chain, users will have less trust in the entire blockchain ecosystem.
@@ I INTRODUCTION
In this paper, we introduce DecTest, a novel decentralized testing framework designed to improve data accuracy. DecTest features a random secret testing mechanism for blockchain oracle systems, revitalizing node monitoring and verification by using a dynamic, anonymized question-verification committee to safeguard data integrity and authenticity. Additionally, we develop an evaluation incentive mechanism that accounts for node reputation scores, offering a fair and transparent assessment of node performance. This approach encourages nodes towards greater honesty and efficiency, ultimately boosting data accuracy.
@@
  • We design a comprehensive incentive mechanism, based on the node’s reputation score for in-depth evaluation, high reputation value will increase the probability of becoming a working node and the rewards will be increased, aiming to motivate oracle nodes and improve the accuracy of data.
@@
  • We creatively propose the Decentralised Testing (DecTest) architecture and conduct a series of experimental evaluations. The experimental results show that the proposed scheme reduces the discrete entropy value of the acquired data for the true value of the data by 61.4% compared to the existing baseline.
@@
The remaining sections of this paper are organized as follows. Section II discusses related work. Section III describes the main elements of the mechanism. Section IV designs the incentive mechanism. Section V presents simulation results and analysis, and Section VI provides the conclusion of this paper.
II Related Work
@@
There has been a large amount of research dedicated to solving the problem of blockchain oracle access to data accuracy, divided into two main directions: reputation-based and voting-based. These solutions provide trusted mechanisms when external data is introduced to the blockchain, giving greater transparency and security to the whole system.
II-A Reputation-Based Oracle
@@
The data provided by oracle is safeguarded from tampering by designing different mechanisms based on reputation. DiOr-SGX [13] uses hardware protection provided by Intel Software Guard Extensions (SGX) to select reputable nodes as leaders for data transfer in a trusted and secure execution environment. Witnet [14] selected nodes to complete RAD tasks and obtain block-producing opportunities based on their reputation values, with higher reputation indicating greater responsibilities. Pasdar et al. [15] developed a livestock blockchain oracle (LBO) that selected nodes based on the oracle’s reputation/performance metrics, addressing the issue of accessing off-chain private sensitive data. Madine et al. [3] designed a reputation system to identify abnormal behavior in oracles, kicking out nodes with low scores from the oracle network. However, using hardware to provide protected data feeds incurs high costs and is relatively difficult to maintain. The use of TLS-type proofs requires the introduction of a third-party proving authority, which has the risk of reducing decentralization.
II-B Voting-Based Oracle
@@
Oracle nodes play roles such as voters and validators to cast votes. ASTRAEA [16] introduced multiple entity users, including validators, submitters, and voters. Submitters delivered challenges into the system, and voters adopted low-risk/low-reward strategies, while validators pursued high-risk/high-reward strategies. Deepthought [17] combined the voting system derived from ASTRAEA with user reputation to reward the most honest users and reduce corruption risks caused by adversarial users or lazy voters. Nelatur et al. [18] proposed a decentralized oracle that uses a crowdsourced voting mechanism to determine the authenticity of queries. Cai et al. [19] presented a scoring scheme based on peer prediction and nonlinear betting rules, aiming to extract subjective data truthfully. Gigli et al. [20] designed oracles that required consensus on specified functions and conducted voting. Vote-based oracle are only rewarded if they agree with the majority result, which can lead to herd behavior. Also, individuals or organizations holding a large number of tokens may try to manipulate the outcome of the vote to serve their interests.
@@
The above design solutions for the oracle focus on data consistency and integrity but do not adequately consider the problem of the oracle being attacked or providing false or deceptive data for selfish motives. Currently, there is almost no mention in existing literature of testing methods to assess the accurate feedback of data from oracle nodes. Oraichain [21] proposed quality testing for AI models, allowing AI providers to earn fees only if their AI models passed these tests, primarily incentivizing AI providers to improve the accuracy of AI models. It is mentioned in [12] that in choosing the most valuable oracle, other oracles will test that oracle randomly, but will be vulnerable to targeted attacks or submission of spoofed data. Simultaneously, due to the public and transparent nature of blockchain, the testing process faces constraints, particularly making it challenging to evaluate the honesty of data feeding by oracle nodes.
III Overview
To improve the security and reliability of the oracle system, we designed a new random secret testing mechanism, DecTest, which mainly contains an incentive mechanism and testing mechanism modules.
III-A DecTest Architecture
@@
We have designed and implemented the DecTest architecture, as shown in Fig. 1. It is mainly divided into four modules: Smart Contract, Oracle Nodes, Data Sources, and Test Cases.
Figure 1: DecTest Architecture.
Smart Contract
@@
Deploy user contracts and oracle contracts on the blockchain, with the former handling blockchain user-initiated data requests. Oracle contracts also include Registration Contracts, Payment Contracts, Reputation Contracts, and Proxy Contracts [22].
Registration Contract. Off-chain nodes register as oracle nodes by pledging virtual currencies, which are securely locked in the payment contract as a strong margin.
@@
Payment Contract. The node completes the task and this deposited amount is converted into a substantial reward, providing a fair reward mechanism for the node’s efforts and contributions.
@@
Reputation Contract. At the end of each round of tasks, the reputation contract calculates and updates the reputation value of the node.
@@
Proxy Contract. Provides an on-chain interface for user contracts and defines the API interface for oracle nodes to fetch data off-chain.
Oracle Nodes
@@
Data Sources
The oracle gets the task and accesses the information through API to off-chain data sources.
Test Cases
Committee members obtain the examples needed for their test tasks from the test case repository [23] for distribution.
III-B Workflow of Random Testing
Fig. 2 illustrates the interactions between various modules.
① When a user needs to get data, he first submits the request $Q$ to the blockchain through a smart contract.
② The nodes are selected based on the Verifiable Random Selection Function (VRF) to form a question-verification committee, and the members generate a test data request $Q$ to be sent to the on-chain smart contract at a certain point in time. (\$III.C.1, \$III.C.2)
③ If the request is for on-chain data, the user contract directly obtains the blockchain data; if the request is for off-chain data, the predictor contract receives and responds to the request, and the blockchain will call the oracle contract after receiving the request from the user contract and passes it the relevant parameters.
④ The oracle node under the chain listens to the request $Q$ or $q$, and we design the node selection scheme based on the reputation value to select the trusted node for the under-chain data feed. (\$III.C.2)
⑤ The selected decentralized oracle goes to off-chain data sources such as servers, the cloud, and the internet to get data based on API.
⑥ The oracle node feeds the data into a smart contract on the blockchain (the node that performs the test data request also feeds the data it fetches onto the chain), and the oracle contract returns the response to the user contract and sends the response to the oracle at the same time.
⑦ The committee members fetch the data delivered by the oracle node from the blockchain and verify that the data fetched by the node is the same as the results of the use case, thus determining whether the oracle node is acting maliciously. (\$III.C.3) The committee feeds the results to the oracle contract to update the reputation value of the oracle node. It also calculates the rewards and punishments. (\$IV)
@@ III-B Workflow of Random Testing
III-C Design of Random Testing
III-C1 System Hypothesis
@@
III-C2 Node Selection Strategy
@@
III-C3 Structure of the question-verification committee
We designed three types of nodes for the question-verification committee, as depicted in Fig. 3 and described below.
Questioner. The questioner’s task is to initiate a query of off-chain data. This selects a test case, publishes the task in the blockchain ecosystem, and then waits for the smart contract to execute the issued request $Q$.
Judge. Judge upon the detection of malicious endeavors, the questioner is summarily excised from the proceedings and subjected to punitive measures, concomitant with the statement of new participants. In this paradigm, the selection of a substitute questioner is predicated on a randomized algorithm, ensuring procedural integrity and fairness. Should a validator discern and report malevolent conduct by the questioner, the judge is tasked with deliberating the veracity of such claims. A determination of guilt results in the accuser assuming the role of the questioner for the ensuing cycle, whereas exoneration entails penalization of the accuser, thus maintaining a self-regulating ecosystem [25].
Validator. The validator is responsible for retrieving the blockchain and validating the data provided by oracle. The results of these validations are then transmitted to the reputation contract, which facilitates the calculation of the reputation score. This mechanism ensures continuous evaluation and validation of off-chain data, thus consolidating the reliability of the system.
@@ III-C3 Structure of the question-verification committee
Figure 3: Question-verification committee structure.
IV Incentive Mechanism
IV-A Reputation of Data-Feeding Oracle
@@
We calculate the reputation score of a node based on its historical cumulative reputation $\mathcal{R}_{acci}$, response time $\mathcal{T}_{i}$, accuracy in completing random test tasks $\mathcal{AC}$, reputation weight $\mathcal{RW}$, and the number of completed test tasks $\mathcal{A}$.
@@
IV-B The Reputation of Committee Members
@@ IV-C Punishment Mechanism
If error $\text{If error instances}>\theta,\text{ then execute}\left\{\begin{array}[]{l}\text{Add to blacklist},\\ \text{Deduct all deposits},\\ \text{Remove node}.\end{array}\right.$
V Experiments and Evaluation
V-A Experiment Settings
@@
V-B Evaluation Results
To highlight the performance of adjusting node selection strategies to encourage oracle honesty based on test results, we will compare our approach with traditional weighted random node selection algorithms, pure random selection algorithms, and Dos network [22]. The discrete entropy value of the data in the acquired data and the real value of the data was reduced by 61.4%. Meanwhile, by adjusting the value of $\alpha$, the $\alpha$ of Our1,2,3 corresponds to [0.01,0.05], [0.1,0.3], [0.4,0.6], respectively, We found small differences in the entropy values of our methods. As shown in Fig. 4, we observe an overall decreasing trend in entropy values for our approach, indicating a gradual concentration of data, while other methods tend to have relatively dispersed data. This suggests that adjusting node selection weights based on test task results significantly reduces the probability of selecting malicious nodes.
In order to study different percentages of malicious nodes, we conducted a series of experiments. In each round of the test task, we recorded the number of nodes with malicious behavior, as shown in Fig. 5, and the results of this experiment clearly demonstrate the process. These malicious nodes try to tamper with the data and we compare the number of truly malicious nodes accurately captured in each round by testing the mechanism under different choices of number of nodes ($n$=30, $n$=50, $n$=80). Fig. 5 represents the number of nodes in the detection system that are truly evil in the case of malicious nodes $s$=10%, 20%, and 30%, in the beginning, the number of malicious nodes detected is high, as the number of tests increases, the number of real malicious nodes is decreasing rapidly, it can be seen that our strategy can inhibit malicious nodes from generating malicious behavior. At the same time, with the selection of 80 decentralized oracles to complete the task, the number of truly malicious nodes is decreasing rapidly with $s$=30malicious nodes, although there is not a big difference between the detection of truly malicious nodes by selecting $n$=30 and $n$=50 nodes in the case of $s$=10malicious nodes.
(a) $s$ = 10%
(b) $s$ = 20%
(c) $s$ = 30%